• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nilanga

    Log entry time 00:03:35 on March29,2005

    Entry number 142658

    keyword=Shift Summary

    16.45: Beam is back, but poor quality: position is off and many trips.

    17.45: Beam acceptable, resume production data.

    22:45: start a new Compton run (needed to reset)

    23:45: noticed SRC lecroy HV is blank, NA rates are fine. Will reset after the current run is finished (#3023).

    continue production data through the shift.